An engineer in a locomotive sees a car stuck on the track at a railroad crossing in front of the train. When the engineer first sees the car, the locomotive is 200 m from the crossing and its speed is 20 m/s. If the engineer’s reaction time is 0.49 s, what should be the magnitude of the minimum deceleration to avoid an accident? Answer in units of m/s 2 .
1
Expert's answer
2020-09-21T08:28:53-0400

Explanations & Calculations


  • The train travels some distance during the engineer's reaction time & if that is S1 then,

"\qquad\qquad
\begin{aligned}
\small S_1 &= \small ut\\
&= \small 20ms^{-1}\times 0.49s\\
&= \small 9.8m
\end{aligned}"

  • Then a distance of "\small (200m-9.8m =) 190.2m" is left for the engineer to stop the train under any acceleration which the train is capable of.
  • Assuming that this deceleration remains constant throughout & applying "\small V^2 =U^2+2as" , deceleration could be found.

"\qquad\qquad
\begin{aligned}
\small s& = \small \frac{V^2-U^2}{2a}\\
\small s &\leq \small 190.2 \,\,\,:\text{to avoid the impact}\\
 \small \frac{0^2-20^2}{2\times a}& \leq \small 190.2 \\
\small a & \geq \small -1.052ms^{-2}\,\,\, : \text{negativity = deceleration}\\
\small |a|_{min} & =\small \bold{1.052ms^{-2}} 
\end{aligned}"

  • Any deceleration of greater value than this stops the train at some distance (of greater safety) before the impact point.